What is the formal charge on carbon of the molecular ion CO32-


(Carbonate ion )? With explaination & PLZ tell me WHAT IS THE MEANING OF FORMAL CHARGE :? PLZ EXPLAIN !!!!!!!


plz reply !!!!

Grade:12

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Priya

a formal charge (FC) is the charge assigned to an atom in a molecule, assuming that electrons in a chemical bond are shared equally between atoms, regardless of relative electronegativity.

The formal charge of any atom in a molecule can be calculated by the following equation:

FC = V - N - {B \over 2}

Where V is the number of valence electrons of the atom in isolation (atom in ground state); N is the number of non-bonding valence electrons on this atom in the molecule; and B is the total number of electrons shared in covalent bonds with other atoms in the molecule.

When determining the correct Lewis structure (or predominant resonance structure) for a molecule, the structure is chosen such that the formal charge (without sign) on each of the atoms is minimized.

Examples:

  • Carbon in methane: FC = 4 - 0 - (8÷2) = 0
  • Nitrogen in NO2-: FC = 5 - 2 - (6÷2) = 0
  • double bonded oxygen in NO2-: FC = 6 - 4 - (4÷2) = 0
  • single bonded oxygen in NO2- FC = 6 - 6 - (2÷2) = -1

for co3-2

V=4, N=0, B=8 (4 from C and 4 from 3 0)

Formal charge on carbon =4-0-8/2= 0

 

All the best.                                                           

AKASH GOYAL

AskiitiansExpert-IIT Delhi

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively :Click here to download the toolbar..